Two collinear harmonic oscillations, having amplitudes 4 and 3 units, respectively, are superposed Calculate the amplitude of the resultant oscillation, their frequencies are equal but they have a phase difference of 90.
1
Expert's answer
2019-05-14T09:47:28-0400
"x_1(t)=4 sin(\omega t)""x_2(t)=3 sin(\omega t + \pi / 2) = 3 cos(\omega t)""x(t)=x_1(t)+x_2(t)=4sin(\omega t)+3cos(\omega t)=A \sdot sin(\omega t +\phi)"

where

"A=\sqrt{3^2+4^2}=5""\phi=arcsin(3/A)=arccos(4/A)"

so, the resultant oscillation:


"x(t)=5 sin(\omega t + arcsin(3/5))"

second option:


"x_2\u200b(t)=3sin(\u03c9t-\u03c0\/2)=-3cos(\u03c9t)""A=\sqrt{(-3)^2+4^2}=5""\phi=arcsin(-3/A)=-arcsin(3/A)"

then, the resultant oscillation:


"x(t)=5 sin(\omega t - arcsin(3/5))"